# SpectralEarth: Training Hyperspectral Foundation Models at Scale

### v1

@@ 6.3 Efficient Fine-tuning
Using a pretrained model can speed up convergence during fine-tuning, hence reducing the computational cost of training deep neural networks. To assess this for our models, we compare the training efficiency of Spec. RN50 models pretrained with DINO relative to training from scratch on the CORINE and CDL benchmarks. Figure 8 displays the evolution of the validation metrics across training epochs. We observe that pretrained models converge more rapidly, particularly in the early stages of training. Notably, fine-tuning matches the 100 epochs performance of training from scratch in $\sim 30$ epochs for EnMAP-CDL. These results highlight the computational advantage of using our pretrained models compared to training from scratch.
@@
Parameter-efficient fine-tuning.
@@
Training with limited labels.
@@
6.4 Cross-Sensor Transferability
To assess the transferability of our pretrained models to other sensors, we evaluate them on the Hyperview and DESIS-CDL datasets. Images in both datasets have been collected by sensors with different spectral characteristics from EnMAP. In Table 4, we evaluate the Spec. RN-50 architecture across different fine-tuning protocols. We observe consistent improvements over training from scratch when using a pretrained model for both datasets. Unsurprisingly, the frozen encoders perform relatively poorly due to the difference in the sensor. However, fine-tuning the spectral adapter is sufficient to close most of the performance gap with fine-tuning. Notably, fine-tuning the adapter only outperforms training from scratch on Hyperview, which can be attributed to overfitting on this dataset.
Evaluation Protocol Hyperview DESIS-CDL
Normalized MSE (%) $\downarrow$ mIoU (%) $\uparrow$
Random Frozen Encoder $96.29$ $48.60$
Training from Scratch $92.09$ $59.85$
Frozen Encoder $92.43$ $51.85$
Fine-tune Adapter $\mathbf{90.08}$ $58.03$
Full Fine-tuning $92.68$ $\mathbf{60.55}$
Table 4: Comparison of SSL versus training from scratch on Hyperview and DESIS-CDL with Spec. RN50 backbone. We use DINO for Hyperview and MoCo-V2 for DESIS-CDL.
6.5 Ablation Studies
Impact of temporal positives.
Temporal positives provide a natural data augmentation for joint-embedding methods and have demonstrated their benefits for multi-spectral imagery [65, 46, 56]. Given that only a subset of SpectralEarth includes temporal positives, it is unclear whether using temporal pairs significantly contributes to the overall performance of the pretrained models. To analyze their effect, we pretrain a Spec. RN50 using DINO with and without temporal pairs. Results in Table 5 confirm that excluding temporal positives degrades performance, with the effect most pronounced in frozen encoder evaluation.
Table 5: Impact of temporal positives. All models are pre-trained on SpectralEarth with MoCo-V2. We report F1 Score for EnMAP-CORINE and mIoU for EnMAP-CDL.
@@ 6.5 Ablation Studies
w/ TP $\mathbf{76.64}$ $\mathbf{77.98}$ $\mathbf{51.53}$ $\mathbf{57.77}$
w/o TP $74.85$ $75.36$ $49.93$ $57.48$
@@
To answer this question, we explore the impact of dataset size by pre-training a Spec. RN50 with DINO on randomly sampled subsets of SpectralEarth with sizes 10K, 25K, 50K, 100K, and 200K locations. We conduct frozen-encoder evaluation on EnMAP-CORINE and EnMAP-CDL. Figure 11 summarizes our findings. We observe a consistent improvement in performance metrics as the scale of the pre-training dataset increases from 10k samples to the full size of SpectralEarth. While the performance on EnMAP-CDL appears to saturate, the F1 Score for EnMAP-CORINE still increases, despite the small size of Spec. RN50. This trend confirms the importance of large-scale datasets for pre-training hyperspectral foundation models.
@@
Importance of the patch size.
Medium-resolution remote sensing imagery requires more fine-grained detail preservation than natural imagery, especially for segmentation tasks. This aspect is sometimes overlooked in the literature, where vision transformers with a patch size of 16x16 or 8x8 are commonly used as backbones to pretrain foundation models to reduce the computational cost  [34, 16, 69, 63]. We argue that a smaller image size with a smaller patch size can lead better trade-offs, particularly for HSI imagery where pixel information is rich and should not be heavily compressed in the embedding space. To analyze the impact of the patch size, we pre-trained Spec. ViT-S using MAE with patch sizes of 4x4, 8x8 and 16x16. Results in Figure 12 show that smaller patches consistently improve performance for segmentation tasks. This is further supported by qualitative results in Figure 13, where models with smaller patches produce more refined segmentation maps. These results are consistent with the performance gap between ViT-based foundation models and ConvNet baselines observed in previous studies [67]. In practice, using additional trainable parameters to extract high-resolution features in combination with a pre-trained ViT can mitigate this issue.
@@
Image Mask 4x4 8x8 16x16
@@
Figure 13: Impact of the patch size in the Spec. ViT-S model on EnMAP-NLCD. From left to right respectively: RGB visualization of the input image, ground truth (GT) mask, and predictions with 4x4, 8x8, and 16x16 patch sizes.
Masking ratio in MAE.
Given our small 4x4 patch size for Spec. ViT models, we expect a higher masking ratio is needed for MAE compared to the usual $75\%$ used in Computer Vision [27]. To confirm this hypothesis, we pre-train Spec. ViT-S using MAE with different masking ratios, from $75\%$ to $99\%$. Given MAE’s poor linear probing results for the classification task, we focus on the segmentation downstream tasks. Results in Figure 14 show that we obtain good representations even with an aggressive $95\%$ masking, which enables faster pre-training as fewer tokens need to go through the encoder.
@@
Table 6: Impact of the spectral adapter in the Spec. RN50 architecture. All models are pre-trained on SpectralEarth with MoCo-V2.
@@
w/o Adapter $72.82$ $77.10$ $47.85$ $55.43$
w/ Adapter $\mathbf{73.97}$ $\mathbf{78.57}$ $\mathbf{51.66}$ $\mathbf{58.10}$
Impact of the spectral adapter.
@@
To assess the impact of the spectral adapter, we compare the classical ResNet-50 with our modified architecture, Spec. RN50. We adjust the first convolutional layer of the ResNet-50 model to accommodate the 202 input bands. Both models are pretrained with MoCo-V2 and evaluated on the EnMAP-CORINE and EnMAP-CDL datasets. The results are summarized in Table 6. Spec. RN50 consistently outperforms the standard ResNet-50 in both frozen encoder evaluation and full fine-tuning. This result confirms the importance of a dedicated spectral feature extractor when dealing with hyperspectral images, particularly for tasks requiring fine-grained spectral discrimination.
7 Conclusion
@@
We introduce SpectralEarth, a large-scale dataset derived from EnMAP imagery as basis for SSL methodologies in the hyperspectral domain. We leveraged SpectralEarth to pre-train foundation models based on popular SSL algorithms. Extensive evaluation on several downstream tasks benchmarked the pre-trained models. Our results demonstrate the effectiveness of models pre-trained on SpectralEarth to improve performance and to reduce computational costs for downstream applications. We believe that the insights derived from this study serve as a basis for future development in SSL for hyperspectral imagery.
@@
As it stands, we encourage further effort to construct additional hyperspectral benchmark datasets. Covering tasks such as unmixing enables a more comprehensive assessment in terms of generalizability of hyperspectral foundation models. In particular, transferability across different hyperspectral sensors remains a challenging problem. Moreover, exploration of a wider range of architectures and SSL algorithms, beyond the Spectral ResNet-50 and ViT models employed in this work, may provide a more comprehensive toolbox for practical deep learning on HSI. Last but not least, exploration and benchmarking multi-sensor SSL methods entails a promising endeavor toward more general geospatial foundation models.
Supplementary Material
8 Implementation Details
8.1 Network Architecture
Spectral Adapter.
The Spectral Adapter consists of three (1D Conv + BN + ReLu) layers. The kernel sizes for the 1D convolutions are 7, 7, and 5, with strides of 5, 5, and 3, respectively. The resulting feature maps have 128 channels. To accommodate different sensors and feed the input into standard 2D backbones, we use a global pooling layer to aggregate the remaining spectral dimensions (if any).
Spec. RN50.
The architecture of Spec. RN50 closely follows the classical ResNet-50 design with two modifications: (a) We replace the stem layers with the spectral adapter without downscaling; (b) The first bottleneck block is adjusted to take an input feature map of 128 channels, as opposed to the ResNet-50’s 64-channel input at this stage.
Spec. ViT-S.
For Spec. ViT-S, we set the number of input channels in the projection layer to 128. A patch size of 4 is chosen based on two considerations: (a) to preserve the rich spectral information in hyperspectral tokens without excessive compression in the patch projection layer, and (b) to preserve fine-grained details spatially while keeping a reasonable compute cost given the small image size of 128x128.
8.2 Training Details
Detailed training configurations and hyperparameters for each downstream task are presented below. For each experiment, a search over the learning rate was performed.
Multi-label Classification
For multi-label classification, we utilized the Binary Cross-Entropy (BCE) loss function. Training was conducted over 100 epochs with a batch size of 128. AdamW optimizer was employed with a cosine annealing scheduler. Data augmentation included random resized crop, horizontal and vertical flipping.
Semantic Segmentation
Training was conducted for 100 epochs on all datasets, using a Cross-Entropy loss. AdamW optimizer was used with a batch size of 64 for EnMAP-NLCD and 32 for EnMAP-CDL and DESIS-CDL. We used a cosine annealing scheduler. Data augmentation included random resized crop, horizontal and vertical flipping.
Regression
We resized all images in the Hyperview dataset to 128x128 pixels. We trained for 200 epochs with a normalized mean squared error (MSE) loss function, as described in Section 5.4. The AdamW optimizer was used along with a cosine annealing scheduler and a batch size of 16. As for the other tasks, we used random resized crop, horizontal and vertical flipping during training.
Figure 15: Sample visualizations of EnMAP tiles used to create the SpectralEarth dataset depicting RGB composites (first row), false color PCA composites (second row), and reflectance values for bands 20 (third row), 120 (fourth row) and 220 (fifth row). These representations highlight the rich information carried in the spectral bands. Notably, each band accentuates distinct features of the terrain.
9 Data Visualizations
In this section, we provide additional data visualizations to get deeper insights into the characteristics of SpectralEarth, as well as hyperspectral EnMAP imagery. Samples of EnMAP tiles, i.e. before division into patches and the filtering of water absorption bands, used to create our dataset are displayed in Figure 15.
(a) Open Water
(b) Developed, Open Space
(c) Developed, Low Intensity
(d) Developed, Medium Intensity
(e) Developed, High Intensity
(f) Barren Land
(g) Deciduous Forest
(h) Evergreen Forest
(i) Mixed Forest
(j) Shrub/Scrub
(k) Grassland/Herbaceous
(l) Pasture/Hay
(m) Cultivated Crops
(n) Woody Wetlands
(o) Emergent Herbaceous Wetlands
Figure 16: Mean spectral reflectance per class in the EnMAP-NLCD dataset.
9.1 Spectral Reflectance Analysis
To gain more insights into the spectral characteristics in our dataset, we present an analysis of per-class spectral reflectance for the NLCD task, visualized in Figure 16.
The spectral reflectance profiles across different land cover classes show diverse characteristics. Classes such as ’Open Water’ and ’Barren Land’ have distinct spectral signatures that are characteristic of their respective land cover types. Conversely, some classes, such as ’Deciduous Forest’ and ’Evergreen Forest’, have profiles that are harder to differentiate from one another. This inter-class similarity poses challenges in classification tasks, highlighting the need for advanced feature extractors capable of capturing such subtle differences.
Furthermore, Figure 16 reveals a high level of intra-class spectral variability, a well-known characteristic of hyperspectral imagery. This variance reflects the diverse nature of each land cover class, influenced by factors such as illumination conditions and seasonal changes. This emphasizes the need for robust models to ensure reliable performance across different geographic regions and environmental conditions.
9.2 Temporal Analysis
Figure 17: SpectralEarth temporal distribution by month.
We analyze the temporal coverage of our dataset in Figure 17. SpectralEarth includes data from all seasons, with a relatively balanced distribution across months. However, certain disparities can be observed, e.g., January is underrepresented while July, August, and March are over-represented. This can be attributed to several factors: (a) Increased cloud cover during winter months, which inherently limits the availability of clear imagery; (b) A satellite outage in December-January 2023 and December-January 2024, which resulted in a temporary interruption of data acquisition; (c) The task scheduling nature of the EnMAP mission. As EnMAP is operationally tasked based on specific use cases, the distribution of data inherently reflects the real-world demands of satellite imagery, leading to a natural variation in seasonal coverage.
9.3 Class Distribution in Downstream Tasks
Figure 18: Class distribution of the EnMAP-CORINE dataset.
Figure 19: Class distribution of the EnMAP-CDL dataset.
Figure 20: Class distribution in the EnMAP-NLCD dataset.
Figure 21: Class distribution in the DESIS-CDL dataset.
EnMAP-CORINE
The EnMAP-CORINE dataset exhibits a diverse range of land cover types. As Figure 18 demonstrates, some classes such as ’Arable Land’ and ’Complex cultivation patterns’ are more frequent, while others like ’Coastal/Inland wetlands’ and ’Beaches, dunes, sand’ are comparatively rare. This class imbalance reflects the natural land cover distribution in Europe.
EnMAP-CDL
The EnMAP-CDL dataset is primarily focused on agricultural crop types. However, following the CDL map, the dataset also includes some non-agricultural classes such as grassland, which is over-represented as shown in Figure 19. Overall, the dataset exhibits a reasonable level of class balance among the different crop types.
EnMAP-NLCD
The EnMAP-NLCD dataset provides a comprehensive view of natural and urban land covers. As depicted in Figure 20, albeit some infrequent classes like ’Developed, High Intensity’, the class distribution is well-balanced.
DESIS-CDL
Similar to the EnMAP-CDL dataset, DESIS-CDL is focused on agricultural classes. Due to the limited availability of DESIS tiles, it was not possible to pair this dataset with EnMAP-CDL. Therefore, the class distribution (see Figure 21) and geographical coverage differ.

### v2

@@ VI-C Efficient Fine-tuning
Using a pretrained model can speed up convergence during fine-tuning, hence reducing the computational cost of training deep neural networks. To assess this for our models, we compare the training efficiency of Spec. RN50 models pretrained with DINO relative to training from scratch on the CORINE and CDL benchmarks. Figure 9 displays the evolution of the validation metrics across training epochs. We observe that pretrained models converge more rapidly, particularly in the early stages of training. Notably, fine-tuning matches the 100 epochs performance of training from scratch in $\sim 10$ epochs for EnMAP-CORINE and EnMAP-CDL. These results highlight the computational advantage of using our pretrained models compared to training from scratch.
@@
Parameter-efficient fine-tuning
@@
Training with limited labels
@@
VI-D Cross-Sensor Transferability
To evaluate the generalizability of SpectralEarth-pretrained models across sensors, we assess their performance on three datasets acquired from sensors not seen during pretraining: EO-1 Hyperion (EO1-CDL), DESIS (DESIS-CDL), and Intuition-1 (Hyperview)  [^6] The data for the hyperview challenge was collected using an airborne sensor that mimics the spectral characteristics of the Intuition-1 mission.. These datasets present different spectral characteristics from EnMAP.
We evaluate Spec. ViT-B and ViT-L models pretrained with MAE and compare them against training from scratch, as well as DOFA and HyperSigma. Our results are summarized in Table IV.
On the CDL-based tasks (EO-1 and DESIS), our models consistently outperform both DOFA and HyperSigma across all protocols. Notably, our frozen encoders outperform HyperSigma EO-1, despite not being pre-trained on EO-1 imagery. This highlights the transferability of features learned from SpectralEarth. Given the relatively small size of the CDL datasets, fine-tuning the adapter performs on par with full fine-tuning and significantly outperforms training from scratch.
On the Hyperview dataset, the best performance is obtained using the frozen Spec. ViT-L encoder. We attribute this to severe overfitting when fine-tuning due to the limited training data and the difficulty of the task. This result demonstrates the practicality of using our pre-trained models as frozen feature extractors.
Model Configuration Hyperview ($MSE_{Norm}$ $\downarrow$) CDL (mIoU $\uparrow$)
Backbone Protocol Intuition-1 EO-1 DESIS
SpecViT-B From Scratch 0.88 $\pm$ 0.01 65.85 $\pm$ 0.78 66.18 $\pm$ 0.50
Frozen 0.85 $\pm$ 0.01 59.02 $\pm$ 0.61 62.22 $\pm$ 0.61
Full FT 0.86 $\pm$ 0.00 66.69 $\pm$ 0.91 68.50 $\pm$ 0.78
Adapter 0.85 $\pm$ 0.00 66.17 $\pm$ 0.58 67.74 $\pm$ 0.29
SpecViT-L From Scratch 0.89 $\pm$ 0.01 65.09 $\pm$ 0.76 66.06 $\pm$ 0.57
Frozen 0.81 $\pm$ 0.02 59.46 $\pm$ 1.02 62.78 $\pm$ 0.31
Full FT 0.87 $\pm$ 0.00 66.34 $\pm$ 0.77 69.05 $\pm$ 0.50
Adapter 0.86 $\pm$ 0.00 66.58 $\pm$ 0.41 68.45 $\pm$ 0.16
DOFA-B Frozen 0.89 $\pm$ 0.00 53.28 $\pm$ 0.82 57.92 $\pm$ 0.37
Full FT 0.87 $\pm$ 0.01 61.09 $\pm$ 0.14 63.43 $\pm$ 0.72
DOFA-L Frozen 0.88 $\pm$ 0.00 50.20 $\pm$ 0.47 58.26 $\pm$ 0.36
Full FT 0.89 $\pm$ 0.01 61.15 $\pm$ 0.68 62.88 $\pm$ 0.34
SpatSigma-B Frozen 0.88 $\pm$ 0.00 57.27 $\pm$ 0.24 59.15 $\pm$ 0.76
Full FT 0.88 $\pm$ 0.00 62.91 $\pm$ 0.16 62.48 $\pm$ 0.29
SpatSigma-L Frozen 0.87 $\pm$ 0.01 51.74 $\pm$ 0.54 56.72 $\pm$ 0.35
Full FT 0.86 $\pm$ 0.01 62.90 $\pm$ 0.67 62.41 $\pm$ 0.35
TABLE IV: Cross-sensor transferability experiments. We pretrain Spec. ViT-B and Spec. ViT-L with MAE on SpectralEarth and evaluate on three datasets from different sensors: Hyperview, EO1-CDL and DESIS-CDL. We compare against training from scratch, DOFA and HyperSigma. Best results are in bold, second best are underlined.
VI-E Ablation Studies
Impact of temporal positives
Temporal positives provide a natural data augmentation for joint-embedding methods and have demonstrated their benefits for multi-spectral imagery [48, 47, 49]. Given that only a subset of SpectralEarth includes temporal positives, it is unclear whether using temporal pairs significantly contributes to the overall performance of the pretrained models. To analyze their effect, we pretrain a Spec. RN50 using DINO with and without temporal pairs. Results in Table V confirm that excluding temporal positives degrades performance, with the effect most pronounced in frozen encoder evaluation.
TABLE V: Impact of temporal positives. A Spec. RN50 model is pre-trained on SpectralEarth with DINO and evaluated on EnMAP-CORINE and EnMAP-CDL.
@@ VI-E Ablation Studies
w/ TP $\mathbf{76.64}$ $\mathbf{77.98}$ $\mathbf{71.15}$ $\mathbf{77.66}$
w/o TP $74.85$ $75.36$ $68.42$ $76.22$
@@
To answer this question, we explore the impact of dataset size by pre-training a Spec. RN50 with DINO on randomly sampled subsets of SpectralEarth with sizes 10K, 25K, 50K, 100K, and 200K locations. We conduct frozen-encoder evaluation on EnMAP-CORINE and EnMAP-CDL. Figure 12 summarizes our findings. We observe a consistent improvement in performance metrics as the scale of the pre-training dataset increases from 10k samples to the full size of SpectralEarth. This highlights the importance of large-scale datasets for hyperspectral self-supervised learning — even with small models such as Spec. RN50.
@@
We also compare SpectralEarth with HySpecNet-11k [14] using the Spec. ViT-B backbone pre-trained with MAE. Table VI reports performance under both frozen encoder and full fine-tuning settings. Pre-training on SpectralEarth consistently outperforms pre-training on HySpecNet-11k, with the largest gains observed in the frozen encoder evaluation. Under complete fine-tuning, the performance gap narrows, as the downstream training offset differences in pretraining.
TABLE VI: Comparison of SpectralEarth and HySpecNet-11k. A Spec. ViT-B is pretrained using MAE and evaluated on EnMAP-CORINE and EnMAP-CDL.
Model EnMAP-CORINE (F1) EnMAP-CDL (mIoU)
Frozen Fine-tune Frozen Fine-tune
HyspecNet-11k $72.39$ $78.20$ $65.73$ $75.20$
SpectralEarth $\mathbf{74.98}$ $\mathbf{79.65}$ $\mathbf{72.42}$ $\mathbf{77.46}$
Importance of the patch size
Medium-resolution remote sensing imagery requires more fine-grained detail preservation than natural imagery, especially for segmentation tasks. This aspect is sometimes overlooked in the literature, where vision transformers with a patch size of $\times16$ or $\times8$ are commonly used as backbones to pretrain foundation models to reduce the computational cost  [73, 50, 70, 16]. We argue that a smaller image size with a smaller patch size can lead better trade-offs, particularly for HSI imagery where pixel information is rich and should not be heavily compressed in the embedding space. To analyze the impact of the patch size, we pre-trained Spec. ViT-S using MAE with patch sizes of $\times4$, $\times8$ and $\times16$. Results in Figure 13 show that smaller patches consistently improve performance for segmentation tasks. This is further supported by qualitative results in Figure 14, where models with smaller patches produce more refined segmentation maps. These results are consistent with the performance gap between ViT-based foundation models and ConvNet baselines observed in previous studies [92]. In practice, using additional trainable parameters to extract high-resolution features in combination with a pre-trained ViT can mitigate this issue.
@@
Image Mask $\times4$ $\times8$ $\times16$
@@
Figure 14: Impact of the patch size in the Spec. ViT-S model on EnMAP-NLCD. From left to right respectively: RGB visualization of the input image, ground truth (GT) mask, and predictions with $\times4$, $\times8$, and $\times16$ patch sizes.
Masking ratio in MAE
Given our small $\times4$ patch size for Spec. ViT models, we expect a higher masking ratio is needed for MAE compared to the usual $75\%$ used in computer vision [22]. To confirm this hypothesis, we pre-train Spec. ViT-S using MAE with different masking ratios, from $75\%$ to $95\%$. Given MAE’s poor linear probing results for the classification task, we focus on the segmentation downstream tasks. Results in Figure 15 show that we obtain good representations even with an aggressive $95\%$ masking, which enables faster pre-training as fewer tokens need to go through the encoder.
@@
TABLE VII: Impact of the spectral adapter in the Spec. RN50 architecture pre-trained on SpectralEarth with MoCo-V2.
@@
w/o Adapter $72.82$ $77.10$ $67.50$ $74.31$
w/ Adapter $\mathbf{73.97}$ $\mathbf{78.57}$ $\mathbf{71.15}$ $\mathbf{77.66}$
Impact of the spectral adapter
@@
To assess the impact of the spectral adapter, we compare the classical ResNet-50 with our modified architecture, Spec. RN50. We adjust the first convolutional layer of the ResNet-50 model to accommodate the 202 input bands. Both models are pretrained with MoCo-V2 and evaluated on the EnMAP-CORINE and EnMAP-CDL datasets. The results are summarized in Table VII. Spec. RN50 consistently outperforms the standard ResNet-50 in both frozen encoder evaluation and full fine-tuning. This result confirms the importance of a dedicated spectral feature extractor when dealing with hyperspectral images, particularly for tasks requiring fine-grained spectral discrimination.
VII Conclusion
@@
We introduce SpectralEarth, a large-scale dataset derived from EnMAP imagery as a basis for SSL methodologies in the hyperspectral domain. We leveraged SpectralEarth to pre-train foundation models based on popular SSL algorithms. Extensive evaluation on several downstream tasks benchmarked the pre-trained models. Our results demonstrate the effectiveness of models pre-trained on SpectralEarth to improve performance and reduce computational costs for downstream applications. We believe that the insights derived from this study serve as a basis for future development in SSL for hyperspectral imagery.
@@
As it stands, we encourage further effort to construct additional hyperspectral benchmark datasets. Covering tasks such as unmixing enables a more comprehensive evaluation of hyperspectral foundation models. In particular, transferability across different hyperspectral sensors remains a challenging problem. Moreover, exploration of a wider range of architectures and SSL algorithms, beyond the Spectral ResNet-50 and ViT models employed in this work, may provide a more comprehensive toolbox for practical deep learning on HSI. Last but not least, exploration and benchmarking of multi-sensor SSL methods is a promising avenue of research toward more general geospatial foundation models.
VIII Acknowledgement
The work of N. Ait Ali Braham was supported by the European Commission through the project “EvoLand” under the Horizon 2020 Research and Innovation program (Grant Agreement No. 101082130). The work of Y. Wang and C. Albrecht was funded by the Helmholtz Association through the Framework of Helmholtz AI, grant ID: ZT-I-PF-5-01 — Local Unit Munich Unit Aeronautics, Space and Transport (MASTr). This work was partially supported by ANR 3IA MIAI@Grenoble Alpes (ANR-19-P3IA-0003), and J. Mairal is supported by ERC grant number 101087696 (APHELEIA project). The compute was supported by the Helmholtz Association’s Initiative and Networking Fund on the HAICORE@FZJ partition. The authors gratefully acknowledge the computational and data resources provided through the joint high-performance data analytics (HPDA) project “terrabyte” of the German Aerospace Center (DLR) and the Leibniz Supercomputing Center (LRZ).